OpenStudy (anonymous):

Solve the system of equations. 3x + 4y = 23 5x + y = 27

OpenStudy (anonymous):

lets multiply the bottom equation by -4 to eliminate y when we add the equations: 3x+4y=23 -20x-4y=-108 add them -17x=-85 x=5 plug back in and y=2

OpenStudy (anonymous):

for solving these type of questions try eliminating a variablbe an solve for the other : multiply the 2nd equation by -4 to eliminate y when we add the equations: 3x+4y=23 -20x-4y=-108 add them -17x=-85 x=5 and so y=2

OpenStudy (anonymous):

multiply up by 1 and down by 4 3x+4y=23 20x+4y=108 then you subtract -17x=-85 x=5 substitute for y

OpenStudy (anonymous):

ok, so we all agree haha

OpenStudy (anonymous):

:D yea we do :D

OpenStudy (anonymous):

So are there 2 answers? y and x?

OpenStudy (anonymous):

yes , since u where asked to solve an equation containing 2 unknows x and y, u will finally find 2 answers .

OpenStudy (anonymous):

Ohh okay.